Operations Report – April 22nd

Crew 315 Operations Report 22-04-2025

SOL: 2

Name of person filing report: Michael Andrews

Non-nominal systems: Main Generator, Spacesuit Charger #7, Perseverance Rover

Notes on non-nominal systems:

1. (Unresolved from yesterday) Oil leak on the main generator, which requires us to use the backup generator at night. This is being worked solely by Mission Support and does not require crew engineering.

2. (Unresolved from yesterday) It was noted that Suit #7 wasn’t charging, but further diagnosis from the Commander has isolated the issue to the #7 charger. Troubleshooting this will take place over the next day.

3. The Perseverance Rover has two minor issues: the parking brake does not fully engage and have much effect, and the ignition block rotates fully if you turn the key too hard by accident.

ROVERS

Spirit rover used: No

Hours: Not Used

Beginning charge: Not Used

Ending charge: Not Used

Currently charging: Not Used

Opportunity rover used:

Hours: 219.0 hours (1.1 hours used)

Beginning charge: 100%

Ending charge: 64%

Currently charging: No

Curiosity rover used: Yes

Hours: 311.8 hours (0.5 hours used)

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: Yes

Hours: 323.8 hours (1.4 hours used)

Beginning charge: 100%

Ending charge: 89%

Currently charging: No

General notes on rovers: None

Summary of Hab operations: Two EVAs were performed today, the first to collect rock samples and the second to take desiccated soil samples. A 100cameras module on camera features was performed with the crew after breakfast. A major repair of the tunnels’ zip ties was also performed (details under RAM notes).

Water Use

Time of measurements (collect data for the same time window): 1930
1) Per formula (see notes): 29.41
2) Smart Home Dashboard (see notes): 25.25

Water (static tank, remaining gallons): 432.37

Static tank pipe heater: Off

Static tank heater: Off

Toilet tank emptied: No

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects

Summary of suits and radios:

· EVA 1: Suits 2, 3 and Radios 2, 3

· EVA 2: Suits 7, 8, 9 and Radios 1, 3, 5

Summary of GreenHab operations: Plants were watered, some basil leaves were harvested for meals. Since the fan doesn’t reliably work due to the main generator being inoperable, the temperature increased quicker than normal and needed to be regulated via opening the GreenHab door.

WATER USE: 10.4 gallons

Heater (On or Off): On

Supplemental light: 1900 – 2200

Harvest (name, weight in grams): Handful of basil leaves (unable to weigh due to faulty scale)

Summary of ScienceDome operations: Began the 3D scanning project with the rock samples recovered on EVA #3. A Rubik’s cube was first used as a simple model to scan, which didn’t work initially. I performed some more research using some downloaded manuals and videos to better understand the software before starting with a first rock sample. The results were much better, yielding a 3D model of a sample after about 2 hours of work. The model was lacking color, but more work will be done to add color to the texture and reduce scanning time. A power study for the EVA Connectivity kit was also performed.

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: (Tools used, work done) put notes here

· Zip Ties: Snips and 44 zip ties were taken from the hab, all to repair the tunnels’ tarps. I’m uncertain if the large amount of zip ties that needed to be replaced was because of the last several days of no engineering support between crew 314 and 315, or if certain zip ties were intentionally not replaced. Approximately half of those zip ties were applied on grommets that didn’t have any nearby zip ties on the ground.

· GreenHab Scale: I hoped that the scale needed a simple AAA battery swap, but it proved to be more complicated. The screen does not turn on even with fresh batteries from the exercise bike. I need to investigate the interior components for any obvious issues.

· Spacesuit #1 Repair: After personally attempting to wear two-piece suit #1, Commander Laude discovered that the neck ring had been installed upside down. I used a screwdriver and pliers to uninstall, flip, and reinstall the neck ring for future use.

· Other hardware used: measuring tape for water usage measurements, and an extension cord and 5 specimen jars were taken from the RAM for ScienceDome projects to be returned at end of mission.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support:

· Is there a way to download the data from the Home Assistant program? That will allow the GreenHab specialist and myself to more accurately record the average temperature and water usage, respectively. Attempting to download the data on the graph view takes us to a new Safari page that cannot be viewed.

· Is there any issue with using a scale in the ScienceDome to weigh harvested plants in the GreenHab? Or must they remain in their respective areas?

Copyright © The Mars Society. All rights reserved. | Main Site